A rock falls from rest off a high cliff. How far has the rock fallen when its speed is 39.2 meters per second? [Neglect friction.]
19.6 m
44.1 m
78.3 m
123 m
-what does neglect friction mean ? i dont quite understand

Answer by scott8148(6628)

friction with the air (or any fluid) will act against the gravitational force that accelerates a falling object
___ this is why skydivers reach a "terminal velocity"; the force of atmospheric friction equals the force of gravity

time = velocity / acceleration

distance = 1/2 * acceleration * (time)^2

substituting ___ distance = 1/2 * (velocity)^2 / acceleration

d = 1/2 * (39.2)^2 / 9.81 = 78.3